OpenStudy (anonymous):

3d + 4 = 2 + 3d - 1/2 SHOW WORK PLEASE

OpenStudy (anonymous):

3ds cancel out so no solution

OpenStudy (anonymous):

3d+4=2+3d-1/2 combine like terms: 3d+4=3d+1.5 move numbers to one side: 3d+2.5=3d answer is no solution